Operations Report – February 6th

Crew 188 Operational Report 06/02/2018

Sol 9

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 15.2h

Solar: SOC 74%

Diesel: 75%

Propane: 61 psi

Ethanol Free Gasoline (5 Gallon containers for ATV): 3 Gallons

Water (trailer): 0 Gallons

Water (static): 170 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130825.4

Toilet tank emptied: Yes

ATVs Used: The three blue ATVs were used today

Oil Added: None

ATV Fuel Used: 2 Gallons

# Hours the ATVs were Used today: 2

Notes on ATVs: It is very difficult to see the fuel level in the ATVs with the suit helmet on. Also, the gas can that had a little fuel in the bottom imploded due to temperature change. There was no leak, but the structural integrity of the can has been compromised and it should be destroyed. I suggest a new location be found to keep the cans out of the sun.

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: Yes

Hours: 19.0h

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 11.4h

Beginning charge: 100%

Ending charge: 99%

Currently charging: No

HabCar: Yes, to get a load of water.

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 9:00

Generator was turned on at 18:04

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Tests were run on the internet connectivity and passed on to Bernard.

Summary of suits and radios: Nominal

Summary of Hab operations: We are running low of vegetables, and there is a concern that our vegetarians will not have enough food for the remainder of the mission.

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.

The “tunnel” had a few sharp edges of the metal at the corners, and a jacket was torn. I’ve taped over the edges so it doesn’t happen again.

Operations Report – February 5th

Crew 188 Operational Report 05/02/2018

Sol 8

Crew reporting: Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 16.9h

Solar: SOC 84%

Diesel: 75%

Propane: 61 psi

Ethanol Free Gasoline (5 Gallon containers for ATV): 5 Gallons

Water (trailer): 0 Gallons

Water (static): 150 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130788.5

Toilet tank emptied: Yes

ATVs Used: Yes, all four were used today.

Oil Added: None

ATV Fuel Used: 3 Gallons

# Hours the ATVs were Used today: 1

Notes on ATVs: It is very difficult to see the fuel level in the ATVs with the suit helmet on.

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: No

Hours: 13.5

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Curiosity rover used: No

Hours: 11.3h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

HabCar: Not used

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 10:36

Generator was turned on at 17:49

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Sub Nominal

Summary of suits and radios: Nominal

Summary of Hab operations: Nominal

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.

Operations Report – February 4th

Crew 188 Operational Report 04/02/2018

Sol 7

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 16.9h

Solar: SOC 80%

Diesel: 75%

Propane: 61 psi

Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons

Water (trailer): 0 Gallons

Water (static): 170 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130739.8

Toilet tank emptied: No

ATVs Used: None

Oil Added: None

ATV Fuel Used: None

# Hours the ATVs were Used today: 0

Notes on ATVs: none

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: No

Hours: 13.5

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Curiosity rover used: No

Hours: 11.3h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

HabCar: Not used

General notes and comments: We are rotating the Rovers with the one plug in we have, and that should be sufficient if we are taking 2 rovers at a time per EVA.

It would seem that the electrical generation isn’t being regulated that well. The diesel and solar power systems will vary between 120V and 140V. No known reason why this is happening.

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 10:36

Generator was turned on at 17:28

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Internet connectivity continues to be a problem even with the data purchased. Devices are having troubles connecting and this is limiting the crew’s ability to monitor the CapCom emails

Summary of suits and radios: The clasp on Suit 1 was bent back to place, it’s as good as new.

Suit 2 was repaired, it was a cold solder joint, as well as the connection not being crimped. I expertly re-attached the connector and the suit is operational

Summary of Hab operations: Nominal

Summary of EVA operations: In the warm weather the fogging issue isn’t as pronounced. We are hoping for warm weather to maintain visual accuracy in the helmets.

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one pack left.

Operations Report – February 3rd

Sol 6

Prepared by: Zac Trolley

Non-Nominal Systems: Greenhab temperature sensor, Suit 2, electrical systems.

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 16.7h
Solar: SOC 84%
Diesel: 76%
Propane: 61 psi
Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons

Water (trailer): 0 Gallons
Water (static): 200 Gallons
Trailer to Static Pump used: No
Water (loft) – Static to Loft Pump used: Yes
Meter 130708.6

Toilet tank emptied: Yes

ATVs Used: None

Oil Added: None

ATV Fuel Used: None
# Hours the ATVs were Used today: 0
Notes on ATVs: none

Deimos rover used: Yes
Hours: Hours: 108.1
Beginning charge: 100%
Ending charge: 98%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 18.0h
Beginning charge: 100%
Ending charge: 91%
Currently charging: No

Opportunity rover used: no
Hours: 13.5
Beginning charge: N/A
Ending charge: N/A
Currently charging: No

Curiosity rover used: Yes
Hours: 11.3h
Beginning charge: 100%
Ending charge: 92%
Currently charging: Yes

HabCar: Not used

General notes and comments: We are rotating the Rovers with the one plugin we have, and that should be sufficient if we are taking 2 rovers at a time per EVA.

It would seem that the electrical generation isn’t being regulated that well. The diesel and solar power systems will vary between 120V and 140V. No known reason why this is happening.

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Summary of internet: Internet connectivity continues to be a problem even with the data purchased. Devices are having troubles connecting and this is limiting the crew’s ability to monitor the CapCom emails

Summary of suits and radios: There is a bent clasp on suit #1 on the helmet collar. It was probably bent during the EVA, the system is still functional.

Suit 2’s air supply is no longer functioning. The diagnostics we performed suggest that the switch isn’t operating properly. We are shelving the EVA suit.

Summary of Hab operations: Nominal

Summary of EVA operations: We are still brainstorming ways to combat the helmet fog. Without proper ventilation under the helmet ring, the facemasks rapidly fog up. It is becoming a safety issue on EVA. The other suit configuration seems to have better airflow.

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: Nominal

Operations Report – February 2nd

Crew 188 Operational Report 02/02/2018

Sol 5

Zac Trolley

Non-Nominal Systems: Propane supply, Microwave, Rover Charging infrastructure, Greenhab temperature sensor, Suit 2, electrical systems.

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 13.5h

Solar: SOC 77%

Diesel: 76%

Propane: 61 psi

Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons

Water (trailer): 0 Gallons

Water (static): 250 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130661.4

Toilet tank emptied: No

ATVs Used: None

Oil Added: None

ATV Fuel Used: None

# Hours the ATVs were Used today: 0

Notes on ATVs: none

Deimos rover used: No

Hours: Hours: 106.9h (Accurate as of two days ago, rover was not on campus during EVA to check hours.)

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes

Hours: 18.0h

Beginning charge: 100%

Ending charge: 91%

Currently charging: No

Opportunity rover used: no

Hours: 13.5

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Curiosity rover used: Yes

Hours: 10.5h

Beginning charge: 100%

Ending charge: 86%

Currently charging: Yes

HabCar: Not used

General notes and comments: The Hab Director mentioned that the Rovers required an extension cord that is shorter than 25 feet. We estimate the cords being used now are 50 foot cables. I was able to find a 25 foot extension cable, meaning we could charge one rover at time based on the above specifications. Can we have the appropriate extension cables sent to the Hab in order to charge the Rovers?

Also, the configuration the rovers are in means that the rovers need to back out before they are used. Backing up in an EVA suit is hazardous since you can’t see behind you in a suit. Please advise if there is another configuration that allows charging with a 25 foot extension cable as well as the ability to drive forward during EVAs.

Summary of internet: Internet connectivity continues to be a problem even with the data purchased. Devices are having troubles connecting and this is limiting the crews ability to monitor the CapCom emails

Summary of suits and radios: There is a bent clasp on suit #1 on the helmet collar. It was probably bent during the EVA, the system is still functional.

Suit 2’s air supply is no longer functioning. The diagnostics we performed suggests that the switch isn’t operating properly. We are shelving the EVA suit.

Summary of Hab operations: Nominal

Summary of EVA operations: We are still brainstorming ways to combat the helmet fog. Without proper ventilation under the helmet ring, the facemasks rapidly fog up. It is becoming a safety issue on EVA. The other suit configuration seems to have better air flow.

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the Hab Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: The propane leak seems to be solved, there was no reports of smell today.

Questions, concerns and requests to Mission Support: The two electrical generation systems are operating on different voltages. The generator is outputting 120V and the solar generator is outputting 140V.

Operations Report – February 1st

Crew 188 Operational Report 01/02/2018
Sol 4
Zac Trolley
Non-Nominal Systems: Propane supply, Microwave, Rover Charging infrastructure, Greenhab temperature sensor. Notes on non-nominal systems: See individual systems at the end of the report. Generator (hours run): 18.8h
Solar: SOC 81%
Diesel: 78%
Propane: 62 psi
Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons Water (trailer): 0 Gallons
Water (static): 300 Gallons
Trailer to Static Pump used: No
Water (loft) – Static to Loft Pump used: Yes
Meter 130607.8
Toilet tank emptied: Yes

ATVs Used: None
Oil Added: None
ATV Fuel Used: None
# Hours the ATVs were Used today: 0
Notes on ATVs: none
Deimos rover used: No
Hours: Hours: 106.9h (Accurate as of two days ago, rover was not on campus during EVA to check hours.)

Beginning charge: N/A
Ending charge: N/A
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: no
Hours: 17.5h
Beginning charge: N/A
Ending charge: N/A
Currently charging: No

Opportunity rover used: no
Hours: 13.5
Beginning charge: N/A
Ending charge: N/A
Currently charging: No

Curiosity rover used: No
Hours: 9.3h
Beginning charge: N/A
Ending charge: N/A
Currently charging: No

HabCar: Not used
General notes and comments: The Hab Director mentioned that the Rovers required an extension cord that is shorter than 25 feet. We estimate the cords being used now are 50 foot cables. I was able to find a 25 foot extension cable, meaning we could charge one rover at time based on the above specifications. Can we have the appropriate extension cables sent to the Hab in order to charge the Rovers?
Also, the configuration the rovers are in means that the rovers need to back out before they are used. Backing up in an EVA suit is hazardous since you can’t see behind you in a suit. Please advise if there is another configuration that allows charging with a 25 foot extension cable as well as the ability to drive forward during EVAs.

Summary of internet: Internet connectivity continues to be a problem even with the data purchased. Devices are having troubles connecting and
Summary of suits and radios: There is a bent clasp on suit #1 on the helmet collar. It was probably bent during the EVA, the system is still functional.
Summary of Hab operations: The microwave isn’t working anymore. The buttons work, the time counts down, but the microwave doesn’t actually produce waves to heat anything.
Summary of EVA operations: We are still brainstorming ways to combat the helmet fog. Without proper ventilation under the helmet ring, the facemasks rapidly fog up. It is becoming a safety issue on EVA.
Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.
Summary of Science Dome operations: The science dome is awaiting a portable heater as per the Hab Director. Summary of RAM operations: Not Operational.
Summary of health and safety issues: The propane detector has arrived, and there is a positive detection of combustible gas in the hab.
We calibrated the sensor outside, and moving the detector in different places in the Hab clearly shows that there are areas where there is a higher concentration of combustible gas than others. The instrument is of low quality and does not show us the ppm, so we have no way to tell what the levels are, and if they are dangerous.
The leak was found using soapy water, and the fitting was tightened by the crew. We will investigate tomorrow and hopefully the leak is fixed. Questions, concerns and requests to Mission Support: All nominal [end]

Operations Report – January 31st

Operational Report
Crew 188: Zac Trolley
31/01/2018

There is a request by the Hab Director for an electrician to investigate a malfunctioning power plug at her residence.
Generator:
The generator shut it’s self off at 4am, the assumed fault was the 12 Volt Battery running low. Luckily the crew was up for the lunar eclipse, and were able to swap the batteries out. The second battery lasted less than 12 hours.
A repair crew came and replaces the alternator on the generator. The oil, fuel, and air filters were also changed. The battery is charging at 14V, so I have high confidence that the issue is solved.

Evac plan
• Generator (hours run): 6.1h
• Generator turned off: 10:33 @ SOC 100%
• Generator turned on: 14:00 @SOC 100%

Fuel Levels:
• Diesel: 79%
• Propane: 62%
• Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons

Water Levels:
• Water (trailer): 0 Gallons
• Water (static): 350 Gallons
• Trailer to Static Pump used: No
• Water (loft) – Static to Loft Pump used: Yes
• Meter @ 20:01 130581.4
• Water Used today: 30.5 Gal
• Toilet tank emptied: No

ATV:
• ATVs were not used today.
• ATVs Used: None
• Oil Added: None
• ATV Fuel Used: None

Rovers:
The Hab Director mentioned that the Rovers required an extension cord that is shorter than 25 feet. We estimate the cords being used now are 50 foot cables. I was able to find a 25 foot extension cable, meaning we could charge one rover at time based on the above specifications. Can we have the appropriate extension cables sent to the Hab in order to charge the Rovers?

Deimos rover used: No
• Hours: 106.9h (Unknown, Director used Rover today, do not have updated numbers)

Spirit rover used: No
• Hours: 17.5h

Opportunity rover used: No
• Hours: 13.5h

Curiosity rover used: No
• Hours: 9.3h

HabCar: Not used

Summary of internet: We purchased the 75 Gig internet token for data in order to support the mission. It is unclear to the crew how the internet tokens operate, and we will monitor the system to ensure we are using is correctly.
Summary of suits and radios: All nominal
Summary of Hab operations: All nominal
Summary of EVA operations: We are still brainstorming ways to combat the helmet fog.
Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.
Summary of Science Dome operations: The science dome is awaiting a portable heater as per the Hab Director.
Summary of RAM operations: Not Operational.
Summary of health and safety issues: There was a report of a brief propane smell today, still awaiting the propane detector to rule out off gassing in the Hab.
Questions, concerns and requests to Mission Support: All nominal

Operations Report – January 30th

Operational Report

Crew 188: Zac Trolley

30/01/2018

Generator:

This morning I was on an EVA, and it was cloudy. Due to the concern from CapCom about the battery bank dropping below 80% I decided to keep the generator running for the morning. Upon my return the generator was no longer running, so I do not have proper metrics for the generator. Shannon suggested that the car battery used to start the generator had died. Upon closer inspection, it the gauge for the battery charger is not working, and It’s suspect that the charger is working at all. I’ve swapped to the backup battery to run the generator tonight.

There as a new battery charger delivered, and it required maintenance to become operational (See attached picture) It is now charging the battery that was hooked up to the generator that died.

· Generator (hours run): Unknown

· Generator turned off: Unknown

· Generator turned on: 17:06 @ SOC 57%

Fuel Levels:

· Diesel: 82%

· Propane: 64%

· Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons

Water Levels:

· Water (trailer): 0 Gallons

· Water (static): 350 Gallons

· Trailer to Static Pump used: No

· Water (loft) – Static to Loft Pump used: Yes

· Meter @ 19:24 -130550.9

· Water Used between 7:00-19:00 37 Gal

· Toilet tank emptied: Yes

ATV:

ATVs were not used today.

· ATVs Used: None

· Oil Added: None

· ATV Fuel Used: None

Rovers:

Spirit and Curiosity were used for today’s EVA. It was noted by the Director that Curiosity was not in the “off” position and not properly charging. This was crew oversight and likely caused by lower visibility in the suit helmets and not being able to see the charging light.

The Hab Director mentioned that the Rovers required an extension cord that is shorter than 25 feet. We estimate the cords being used now are 50 foot cables. I was able to find a 25 foot extension cable, meaning we could charge one rover at time based on the above specifications. Can we have the appropriate extension cables sent to the Hab in order to charge the Rovers?

Deimos rover used: No

· Hours: 106.9h

Spirit rover used: Yes

· Hours: 17.5h

· Beginning charge: 63%

Opportunity rover used: No

· Hours: 13.5h

· Beginning charge: 100%

Curiosity rover used: Yes

· Hours: 9.3h

· Beginning charge: 56%

HabCar: Not used

Summary of internet: We will be purchasing 5 gigs of data per day in order to meet our Sim objectives and be able to increase the visibility of this mission.

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of EVA operations: The helmets fogging continue to be an issue. We are going to bring the helmets into the open area in an effort warm up the Lexan before we go on an EVA. Please advise if there are other tricks and tip we can use to lower the fogging issue. It’s becoming a safety issue as visibility is now an unknow variable.

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the Hab Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: There was no report of a propane smell today, still awaiting the propane detector to rule out off gassing in the Hab.

Questions, concerns and requests to Mission Support: A crew member had previous version of the Hab hand book (2003, 2016) and those manuals mentioned higher functionality in the Hab in some areas (HALpr, weather station, CapCom and local chat room, etc) I am interested in bringing similar capability and more to the Hab; who is a good contact Earth side to discuss these improvements while I’m at the Hab and can investigate?

Operations Report – January 29th

Prepared by: Zac Trolley, Operations Manager

Non-nominal systems: All systems are operating in their handover condition
Notes on non-nominal systems: None.

Generator (hours run): 17.1h
Generator turned off: 10:23
Generator turned on: 17:38

Solar
SOC 7:37 93%
SOC 17:38 85%

Diesel: 83%
Propane: Information to be provided by the Director
Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons
Water (trailer): 0 Gallons
Water (static): 400 Gallons
Trailer to Static Pump used: No
Water (loft) – Static to Loft Pump used: Yes
Water Meter: @17:00 130497.5
Toilet tank emptied: No
ATVs Used: HONDA, 350(1), 350(2), 350(3)
All ATV’s used for a short EVA to the marbles. No fuel was added.
Oil Added: added to ATV #2
ATV Fuel Used: None
# Hours the ATVs were used today: 0.5
Notes on ATVs: None

Deimos rover used: Yes
Hours: 106.9h
Beginning charge: 100%
Ending charge: 100%
Currently charging: no

Spirit rover used: Yes
Hours: 16.4h
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: Yes
Hours: 13.5h
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Curiosity rover used: Yes
Hours: 7.9h
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No
General notes and comments: None
Summary of internet: All nominal
Summary of suits and radios: We had an issue where one of the Radios got stuck on transmit on channel 12, causing a comms blackout. This happened with everyone in the Hab, getting ready for an EVA. The Radio was found and the problem was solved before the EVA or anyone left the Hab.
Summary of Hab operations: All nominal
Summary of EVA operations: Had an issue with the facemasks fogging up. Several people had issues with visibility, one person needed to be driven back due to visibility. We will investigate an airflow solution to provide more anti fogging ability.
Summary of GreenHab operations: All nominal
Summary of Science Dome operations: All nominal.
Summary of RAM operations: Not Operational.
Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: Quality data comes from the consistent method of collection of data. To that end, I have a few questions about the data being collected.

For the state of charge, what is that data being used for, and when should it be taken? Without the SOC being tied to the time of day and amount of sun, it doesn’t have much value. (unless I’m missing something)

The water number has a similar issue. Knowing what time of day it was taken, and taking it at the same time every day provides more useful data. I’ve been writing down the water flow several times a day to track our own use. What would you like the information to accomplish?

Last, the information on the rovers won’t be accurate, meaning if I plugged them in, the battery percentage won’t be accurate at the time of this report. Is that to keep an eye on the battery function? If so, I will only transmit information that is relevant to that investigation.

Operations Report – January 28th

Operational Report

Crew 188: Zac Trolley

28/01/2018

Non-nominal systems: None

Notes on non-nominal systems:

Generator (hours run): 14h

Generator turned off: 9:50

Generator turned on: 17:45

Solar— SOC

@9:50 80%

@13:00 100%

@17:45 82%

Diesel: 85%

Propane: 75%

Ethanol Free Gasoline (5 Gallon containers for ATV): 10 Gallons

Water (trailer): 0 Gallons

Water (static): 400 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: @19:00 130469.3

Toilet tank emptied: No

ATVs Used: HONDA, 350(1), 350(2), 350(3)

Oil Added? No

ATV Fuel Used: Very little in learning how to fuel the ATVs

# Hours the ATVs were used today: 0.5h Training

Notes on ATVs: Took the ATVs on a training run. The hours run have not been updated from yesterday.

Deimos rover used:

Hours: 102.3 (crew) + Director discretional hours

Beginning charge: 100%

Ending charge: –

Currently charging: Maybe

Sojourner rover used: Assigned to director only.

Hours: Director discretional hours

Beginning charge:

Ending charge: –

Currently charging: Maybe

Spirit rover used: Yes

Hours: 15.1h

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

Opportunity rover used: Yes

Hours: 12.4h

Beginning charge: 100%

Ending charge:

Currently charging: No

Curiosity rover used: Yes

Hours: 6.6h

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

HabCar used and why, where? Yes, Water run.

General notes and comments: There is a faint smell of propane in the Hab, we believe it may be from the heater on the lower floor when we tried to turn it on. There was also a minor smell of fuel gas during the day.

Summary of internet: All nominal

Summary of suits and radios: Not checked

Summary of Hab operations: All Nominal

Summary of EVA operations: N/A

Summary of GreenHab operations: All nominal

Summary of Science Dome operations: All nominal.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: N/A